OpenStudy (anonymous):

HELP!!?? What percentage of 128 is 96?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

96 parts in 128 converted to x parts in 100

OpenStudy (anonymous):

what does that mean

OpenStudy (anonymous):

128% of 96=122.88

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

I did the math I got how I did it right here you want me to give you it

OpenStudy (anonymous):

but that not the answer

OpenStudy (anonymous):

i need someone to explain how to do the problem

OpenStudy (anonymous):

ohh okay ill tell you how to do it

OpenStudy (anonymous):

To get the solution, we are looking for, we need to point out what we know. 1. We assume, that the number 96 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 96 is 100%, so we can write it down as 96=100%. 4. We know, that x is 128% of the output value, so we can write it down as x=128%. 5. Now we have two simple equations: 1) 96=100% 2) x=128% where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 96/x=100%/128% 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. 7. Solution for what is 128% of 96 96/x=100/128 (96/x)*x=(100/128)*x - we multiply both sides of the equation by x 96=0.78125*x - we divide both sides of the equation by (0.78125) to get x 96/0.78125=x 122.88=x x=122.88 now we have: 128% of 96=122.88

OpenStudy (anonymous):

please A. 75% B. 80% C. 70% D. 65%

OpenStudy (anonymous):

ohh okay hold on

OpenStudy (anonymous):

133.33333333333 okay now it is your turn to do the math

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

\[\frac{96}{128}=0.75=75 \% \]

OpenStudy (anonymous):

how did you get 0.75?

OpenStudy (anonymous):

you have to divide it

OpenStudy (anonymous):

yes it equals 133.3333 thats not the answer

OpenStudy (anonymous):

133.3333 is not the awnser

OpenStudy (anonymous):

Is 64 50% of 128 ? Is 128 100% of 128? Is 96 75% pf 128? What percentage of 128 is 96?

OpenStudy (anonymous):

basiclly dividing